OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? x + y = 8 x - y = 10

OpenStudy (anonymous):

Two equations in the system: 1 x + 1 y = 8 1 x - 1 y = 10 Matrix notation: \[\left[\begin{matrix}1 & 1 \\ 1 & -1\end{matrix} | \begin{matrix}8 \\ 10\end{matrix} \right]\] The system determinant takes the coefficients A to the left (the x,y values) and calculates a value for this square matrix. \[\left| \left[\begin{matrix}a & b \\ d & e\end{matrix}\right] \right| = ae - bd\]